In the month of June, Jose Hebert’s Beauty Salon gave 3,900 haircuts, shampoos, and permanents at an average price of $30. During the month, fixed costs were $16,800 and variable costs were 75% of sales.

Determine the contribution margin in dollars, per unit and as a ratio. (Round contribution margin per unit and contribution margin ratio to 2 decimal places, e.g. 5.25 & 10.50.)

Contribution margin

$

Contribution margin per unit

$

Contribution margin ratio %

eTextbook and Media

Using the contribution margin technique, compute the break-even point in dollars and in units. (Round answers to 0 decimal places, e.g. 1,225.)

Break-even point

$

Break-even point units

eTextbook and Media

Compute the margin of safety in dollars and as a ratio. (Round answers to 0 decimal places, e.g. 1,225.)

Margin of safety

$

Margin of safety ratio %
Answer #1

Contribution margin = (30-30*75%)*3900 = $29250

Contribution margin per unit = 30 -30*75% = $7.5 per unit

Contribution margin ratio = 7.5/30 = 25%

Break even point = Fixed cost/ contribution margin

= 16800/0.25 = $66000

Break even point = 16800/7.5 = 2240 units

Margin of safety = Sales - break even sales

= 30*3900 - 66000 = $51000

Margin of safety ratio = 51000 / 117000 = 43.59%